A-Zone Media sells two models of e-readers. The budgeted price per unit for the wireless model is $205 and the budgeted price per unit for the wireless and cellular model is $442. The master budget called for sales of 11,300 wireless models and 3,150 wireless and cellular models during the current year. Actual results showed sales of 8,800 wireless models, with a price of $265 per unit, and 4,750 wireless and cellular models, with a price of $660 per unit. The standard variable cost per unit is $93 for a wireless model and $225 for a wireless and cellular model.


 


Required:


a. Compute the activity variance for these data. (Do not round intermediate calculations. Indicate the effect of each variance by selecting "F" for favorable, or "U" for unfavorable. If there is no effect, do not select either option.)


 


b. Compute the mix and quantity variance for these data. (Do not round intermediate calculations. Indicate the effect of each variance by selecting "F" for favorable, or "U" for unfavorable. If there is no effect, do not select either option.)
Answer #1

A) 1) Flexible Budget AQ × (SP – SV)

=8800*($205-$93)+4750*($442-$225)

=8800*$112+4750*$217

=2016350

2) ASQ × (SP – SV)  

=13550*(11300/14450)*$205-$93

=2172126.723

3) 13550*(3150/14450)*$442-$225

=1305357.353

then last 2) and 3) total

=2172126.723+1305357.353

=3477484.076

=2016350-3477484.076

Mix Variance= -1461134.08 (U)

B) ASQ × (SP – SV)=$3477484.076

C) Master Budget-

=11300*($205-$93)+3150*($442-$225)

=1949150

Mix Variance= -$1461134.08 (U)

Quantity Variance= $1528334. 08(F)

activity variance=$-67200(U)